# Heterogeneous network estimation for single-cell transcriptomic data via a joint regularized deep neural network

### v1

@@ Abstract
Network estimation has been a critical component of single-cell transcriptomic data analysis, which can provide crucial insights into the complex interplay among genes, facilitating uncovering the biological basis of human life at single-cell resolution. Despite notable achievements, existing methodologies often falter in their practicality, primarily due to their narrow focus on simplistic linear relationships and inadequate handling of cellular heterogeneity. To bridge these gaps, we propose a joint regularized deep neural network method incorporating a Mahalanobis distance-based K-means clustering (JRDNN-KM) to estimate multiple networks for various cell subgroups simultaneously, accounting for both unknown cellular heterogeneity and zero-inflation and, more importantly, complex nonlinear relationships among genes. We innovatively introduce a selection layer for network construction and develop homogeneous and heterogeneous hidden layers to accommodate commonality and specificity across multiple networks. Through simulations and applications to real single-cell transcriptomic data for multiple tissues and species, we show that JRDNN-KM constructs networks with more accuracy and biological interpretability and, meanwhile, identifies more accurate cell subgroups compared to the state-of-the-art methods in the literature. Building on the network construction, we further find hub genes with important biological implications and modules with statistical enrichment of biological processes.
@@ 1 Introduction
Exploring biological networks has become essential for unraveling the complex molecular systems that sustain life. These networks, involving sophisticated interactions among genes, proteins, and other cellular components, are crucial for grasping how cells behave, regulate themselves, and how diseases arise (Costanzo et al., 2019; Zheng & Tang, 2024). In recent years, the explosion of biological data generated through high-throughput omics technologies has propelled the importance of network estimation research to unprecedented heights. Among the existing techniques, the Gaussian Graphical Models (GGMs) are perhaps the most popular. In GGMs, a sparse precision matrix (the inverse of the correlation matrix) is estimated to infer the network (Zhang & Zou, 2014; Halama et al., 2024), describing the conditional dependencies between cellular components given the rests. The estimation procedure can be simplified into a set of sparse node-wise linear regression problems with simpler optimization. This conditional construction strategy takes a more systematic perspective than the simple unconditional construction (where “the other components” are ignored, such as the Pearson correlation-based methods), potentially leading to more biologically implicated results.
@@
A few other studies conduct a further step and additionally take into account the cellular heterogeneity and construct multiple networks for different cell subgroups, where the interactions, coordination, and other relationships among genes may vary. Examples include the Bayesian latent Gaussian graph mixture model (BLGGM) (Wu & Luo, 2022), which can handle the unknown cellular heterogeneity and zero inflation simultaneously. Aside from the heterogeneity of the networks, the potentially shared common structures within cell subgroups have received great attention, as cells belonging to distinct subgroups are frequently derived from the same tissue. Motivated by some studies originally developed for bulk genomics data, which are based on the fused and group Lasso penalties, a few joint network estimation methods have been proposed for effectively accommodating both the heterogeneity and homogeneity among cell subgroups, as well as the unique characteristics of single-cell transcriptomic data. Examples include the joint Gaussian copula graphical model (Wu et al., 2020), GGM incorporating a Bayesian zero-inflated Poisson (ZIP) model (Dong et al., 2023), and kernelized multiview signed graph learning (Karaaslanli et al., 2023).
@@
Despite considerable successes, the aforementioned GGM-based methods only focus on linear dependencies, limiting their ability to capture complex nonlinear relationships among genes that are involved in almost all biological processes. To fill this gap, a few studies take advantage of the nonparametric or distribution-free technique to detect the potential nonlinear relationships, such as the tree-based GENIE3 method (Huynh-Thu et al., 2010), the nonparametric test-based locCSN method (Wang et al., 2021), and the distribution-free-based CS-CORE method (Su et al., 2023). However, these methods mostly adopt the unconditional construction strategy and were originally developed for a single network analysis and depend on the prior known cell subgroup information, which may still lose effectiveness under the practical scenarios with usually complex conditional dependence and unknown cellular heterogeneity.
@@
In response to these limitations, this work develops a novel joint regularized deep neural network incorporating a Mahalanobis distance-based K-means clustering (JRDNN-KM) for single-cell transcriptomic data, achieving network estimation and cell subgroup identification simultaneously. A workflow of JRDNN-KM is presented in Figure 1. Significantly advancing from the published studies, by exploiting the neural network’s capacity for flexible function approximation, JRDNN-KM can capture various nonlinear dependencies and interactions among genes in different cell subgroups while concurrently handling zero-inflation. By employing sparse and similarity regularization techniques, we ensure that the networks also share some common structures, enabling a more comprehensive joint exploration of different cell subgroups. The inferred networks are further incorporated into the Mahalanobis distance-based K-means clustering for detecting unknown cell subgroups, with the advantage of accommodating both the mean and covariance heterogeneous among cells. We demonstrate the performance of JRDNN-KM through extensive simulations and real single-cell transcriptomic datasets from humans and mice. JRDNN-KM detects biologically sensible cell subgroups, hub genes, and modules, significantly contributing to revealing the biological mechanisms driving cellular processes.
@@
Figure 1: Workflow of JRDNN-KM. (A) Input: normalized single-cell transcriptome data $\boldsymbol{x}_{i}$’s and initialized cell subgroup memberships $C_{i}$’s. (B) JRDNN-KM: iterations between JRDNN and Mahalanobis distance-based K-means clustering. (C) Output: $K$ networks for the $K$ subgroups, with both common and specific edges, and estimated cell subgroup memberships. (D) Network structure of the proposed JRDNN, including the selection layers and combinations of homogeneous and heterogeneous hidden layers, with a zero-inflated-based loss incorporating sparse and similarity regularization terms.
2 Methods
@@
Suppose there are $n$ independent cells from $K$ subgroups. For the $i$th cell, denote $C_{i}\in\{1,\ldots,K\}$ as the subgroup assignment and $\boldsymbol{y}_{i}=\left(y_{i1},\ldots,y_{ip}\right)^{\top}$ as the $p$-dimensional vector of the observed count transcriptomic data. In practice, $\boldsymbol{y}_{i}$ has a high sparsity with zero-inflation due to the dropout event. Following the published studies (Booeshaghi & Pachter, 2021; Hafemeister & Satija, 2019), we conduct normalization using the R package Seurat and $\log(1+x)$ transformation on $\boldsymbol{y}_{i}$’s to accommodate the library size and count nature and denote the processed continuous vector as $\boldsymbol{x}_{i}=\left(x_{i1},\ldots,x_{ip}\right)^{\top}$.
@@ 2.1 Network estimation based on a joint regularized deep neural network
First, assume that the subgroup assignment is known, and for each subgroup, each gene is standardized to have a zero mean and unit variance. For accommodating the complex interplay among genes and zero inflation, when $C_{i}=k$, we consider a zero-inflated conditional Gaussian distribution as:
$x_{ij}|\boldsymbol{x}_{i,\backslash j};C_{i}=k\sim$ $\mathcal{N}\left(f_{kj}(\boldsymbol{w}_{k,\backslash j}*\boldsymbol{x}_{i,\backslash j}),1\right)\pi_{j}+\delta_{0}(x_{ij})(1-\pi_{j}),$ (1)
where $\boldsymbol{x}_{i,\backslash j}=\left(x_{i,1},\ldots,x_{i,j-1},x_{i,j+1},\ldots,x_{i,p}\right)^{\top}$, $\boldsymbol{w}_{k,j}=(w_{k,j1},\ldots,w_{k,j(j-1)},w_{k,j(j+1)},\ldots,w_{k,jp})^{\top}$ is a sparse weight vector for $\boldsymbol{x}_{i,\backslash j}$, * is the element-wise product, $f_{kj}(\boldsymbol{w}_{k,j}*\boldsymbol{x}_{i,\backslash j})$ is the mean parameter of Gaussian distribution with $f_{kj}(\cdot)$ being an arbitrarily nonparametric function, $\delta_{0}(\cdot)$ is a Dirac probability measure with a point mass at zero, and $0\leq\pi_{j}\leq 1$ is the probability that the $j$th gene does not express zero caused by dropouts.
Based on Model (1), when $C_{i}=k$ and $x_{ij}$ does not express zero, we have $x_{ij}=f_{kj}(\boldsymbol{w}_{k,j}*\boldsymbol{x}_{i,\backslash j})+\varepsilon_{kj}$ with $\varepsilon_{kj}\sim\mathcal{N}\left(0,1\right)$, where $f_{kj}(\cdot)$ describes the nonlinear relationships between the $j$th gene and the others for the $k$th cell subgroup. Here, we introduce the sparse weight vector $\boldsymbol{w}_{k,j}$ to accommodate the sparse connections among genes and further construct the network. Specifically, if $w_{k,jl}\neq 0$ or $w_{k,lj}\neq 0$, the $j$th and $l$th genes have nonlinear conditional dependence and will be connected in the network. Model (1) can be treated as a nonlinear extension of the GGM-based node-wise regression (Cai et al., 2016; Wang et al., 2019).
For estimating $f_{kj}(\cdot)$ and the unknown parameters $\boldsymbol{w}_{k,j}$, we propose a joint regularized deep neural network (JRDNN) as shown in Figure 1(D). Specifically, for each $j$ and $k$, the function represented by JRDNN can be written as (we omit the dependence on $j$ to simplify notation):
@@
Here, $\boldsymbol{z}_{i}\in\mathbb{R}^{(p-1)\times 1}$ is an input vector, which is $\boldsymbol{x}_{i,\backslash j}$ in our network estimation procedure. We introduce a selection layer for $\boldsymbol{w}_{k}$ and consider $M$ hidden layers. Specifically, in (4), we consider two sub-fully connected neural networks, consisting of heterogeneous neurons and homogeneous neurons, respectively, where heterogeneous neurons and homogeneous neurons are connected with themselves but not connected with each other. Specifically, for the $m$th hidden layer with $m=1,\cdots,M$, we consider $d_{k}^{(m)}$ heterogeneous neurons specific to the $k$th subgroup and $d_{0}^{(m)}$ homogeneous neurons shared by all $K$ subgroups, and $T_{k}^{(m)}(\boldsymbol{u})=\boldsymbol{\Delta}^{(m)}_{k}\boldsymbol{u}+\boldsymbol{b}_{k}^{(m)}$ and $T_{0}^{(m)}(\boldsymbol{u})=\boldsymbol{\Delta}^{(m)}_{0}\boldsymbol{u}+\boldsymbol{b}_{0}^{(m)}$ are affine transformations involving unknown parameters $\boldsymbol{\Delta}^{(m)}_{k}\in\mathbb{R}^{d_{k}^{(m)}\times d_{k}^{(m-1)}}$ and $\boldsymbol{b}_{k}^{(m)}\in\mathbb{R}^{d_{k}^{(m)}\times 1}$ and $\boldsymbol{\Delta}^{(m)}_{0}\in\mathbb{R}^{d_{0}^{(m)}\times d_{0}^{(m-1)}}$ and $\boldsymbol{b}_{0}^{(m)}\in\mathbb{R}^{d_{0}^{(m)}\times 1}$, respectively. $h(\cdot)$ is the activation function, which can be rectified linear unit (ReLU), sigmoid, tanh, and some others. In addition, an output layer with $T_{k}^{(M+1)}(\boldsymbol{u})=\boldsymbol{\Delta}^{(M+1)}_{k}\boldsymbol{u}+\boldsymbol{b}_{k}^{(M+1)}$ is introduced for the output $x_{ij}$, with $\boldsymbol{\Delta}^{(M+1)}_{k}\in\mathbb{R}^{1\times\left(d_{k}^{(M)}+d_{0}^{(M)}\right)}$ and $\boldsymbol{b}_{k}^{(M+1)}\in\mathbb{R}^{1\times 1}$, which integrates both the heterogeneous and homogeneous neurons. Denote $\boldsymbol{U}_{k}$ as a vector consisting of all parameters in the $M$ hidden layers and output layer.
@@
$l_{pn}(\boldsymbol{X};\Phi)$ $=\sum_{j=1}^{p}\left\{-\sum_{i=1}^{n}\sum_{k=1}^{K}\tau_{ik}l_{n}(\boldsymbol{x}_{i};\boldsymbol{U}_{kj},\boldsymbol{w}_{k,j},\pi_{j})+\lambda_{1}\sum_{l\neq j}\sum_{k=1}^{K}|w_{k,jl}|\right.$
@@
with $l_{n}(\boldsymbol{x}_{i};\boldsymbol{U}_{kj},\boldsymbol{w}_{k,j},\pi_{j})=\delta_{ij}\left[\log\pi_{j}-\frac{1}{2}\left\{\log(2\pi_{j})+\left(x_{ij}-g(\boldsymbol{x}_{i,\backslash j};\boldsymbol{U}_{kj},\boldsymbol{w}_{k,j})\right)^{2}\right\}\right]+(1-\delta_{ij})\log(1-\pi_{j}),$ where $\boldsymbol{X}$ is a $n\times p$ matrix consisting of $\boldsymbol{x}_{1},\cdots,\boldsymbol{x}_{n}$, $\Phi$ denotes all unknown parameters, $1(\cdot)$ is an indicator function, $\tau_{ik}=1(C_{i}=k)$, $\delta_{ij}=1(x_{ij}\neq 0)$, and $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are three non-negative tuning parameters.
In (2.1), the first term is the negative log-likelihood function, which models the fit of our data. The second term, an $\ell_{1}$ penalty, encourages sparsity in the selection layer parameters $w_{k,jl}$’s to facilitate sparse network estimation. The third and fourth terms are designed to accommodate the commonality among different cell subgroups. Specifically, for any pairs of $j$ and $l$, the third term promotes $1(w_{k,jl}\neq 0)=1(w_{k^{\prime},jl}\neq 0)$, resulting in that $w_{k,jl}$ and $w_{k^{\prime},jl}$ tend to be zero or nonzero simultaneously. Thus, the $K$ subgroups potentially share some common edges. We further introduce the fourth term to accommodate the similarity in connection strength of the edges. Specifically, the fourth term imposes the F-norm on the weights $\boldsymbol{\Delta}^{(1)}_{k}$ of heterogeneous neurons in the first layer, promoting the elements of $\boldsymbol{\Delta}^{(1)}_{k}$ shrunk towards zero simultaneously and thus pruning all heterogeneous neurons in the $M$ hidden layers. With (2.1), the constructed networks of different cell subgroups not only reflect heterogeneity but also exhibit common connections with similar strength.

### v2

# Heterogeneous gene network estimation for single-cell transcriptomic data via a joint regularized deep neural network

@@ Abstract
Estimation of intracellular gene networks has been a critical component of single-cell transcriptomic data analysis, which can provide crucial insights into the complex interplay between genes, facilitating the discovery of the biological basis of human life at single-cell resolution. Despite notable achievements, existing methodologies often falter in their practicality, primarily due to their narrow focus on simplistic linear relationships and inadequate handling of cellular heterogeneity. To bridge these gaps, we propose a joint regularized deep neural network method incorporating Mahalanobis distance-based K-means clustering (JRDNN-KM) to estimate multiple networks for various cell subgroups simultaneously, accounting for both unknown cellular heterogeneity and zero inflation, and, more importantly, complex nonlinear relationships among genes. We introduce an innovative selection layer for network construction, along with hidden layers that include both shared and subgroup-specific neurons, to capture common patterns and subgroup-specific variations across networks. Applied to real single-cell transcriptomic data from multiple tissues and species, JRDNN-KM demonstrates higher accuracy and biological interpretability in network estimation, and more accurately identifies cell subgroups compared to current state-of-the-art methods. Building on network construction, we further find hub genes with important biological implications and modules with statistical enrichment of biological processes.
@@ 1 Introduction
Exploring intracellular gene networks has become essential for unraveling the complex molecular systems that sustain life. These networks, involving sophisticated interactions among genes, are crucial for grasping how cells behave, regulate themselves, and how diseases arise (Costanzo et al., 2019). In recent years, the explosion of high-throughput omics data has propelled the importance of network estimation research to unprecedented heights. Among the existing techniques, the Gaussian Graphical Models (GGMs) are perhaps the most popular. In GGMs, a sparse precision matrix (the inverse of the covariance matrix) is estimated to infer the network (Zhang and Zou, 2014; Halama et al., 2024), describing the conditional dependencies between genes given the rests. The estimation procedure can be simplified into a set of sparse node-wise linear regression problems with simpler optimization. Conditional dependence networks outperform correlation-based methods (where “the other components” are ignored) by avoiding misleading marginal dependencies and providing mechanistic insights. Crucially, they enable more accurate reconstruction of gene networks by distinguishing direct molecular interactions from indirect associations, whereas correlation-based approaches may conflate these relationships due to unaccounted system-level dependencies.
@@
A few other studies conduct a further step and additionally take into account the cellular heterogeneity and construct multiple networks for different cell subgroups, where the interactions, coordination, and other relationships among genes may vary. Examples include the Bayesian latent Gaussian graph mixture model (BLGGM) (Wu and Luo, 2022), which can handle the unknown cellular heterogeneity and zero inflation simultaneously. Aside from the heterogeneity of the networks, the potentially shared common structures within cell subgroups have received great attention, as cells belonging to distinct subgroups are frequently derived from the same tissue. Motivated by some studies originally developed for bulk genomics data, which are based on the fused and group Lasso penalties, a few joint network estimation methods have been proposed for effectively accommodating both the heterogeneity and homogeneity among cell subgroups, as well as the unique characteristics of single-cell transcriptomic data. Examples include the joint Gaussian copula graphical model (Wu et al., 2020), GGM incorporating a Bayesian zero-inflated Poisson (ZIP) model (Dong et al., 2023), and kernelized multiview signed graph learning (Karaaslanli et al., 2023).
@@
Despite considerable successes, the aforementioned GGM-based methods only focus on linear dependencies, limiting their ability to capture complex nonlinear relationships among genes that are involved in almost all biological processes. To fill this gap, a few studies take advantage of the nonparametric or distribution-free technique to detect the potential nonlinear relationships, such as the tree-based GENIE3 method (Huynh-Thu et al., 2010), the nonparametric test-based locCSN method (Wang et al., 2021), and the distribution-free-based CS-CORE method (Su et al., 2023). However, these methods mostly adopt the unconditional construction strategy and were originally developed for single-network analysis and depend on the prior known cell subgroup information, which may still lose effectiveness under the practical scenarios with usually complex conditional dependence and unknown cellular heterogeneity.
@@
In response to these limitations, this work develops a novel joint regularized deep neural network incorporating a Mahalanobis distance-based K-means clustering (JRDNN-KM) for single-cell transcriptomic data, achieving intracellular gene network estimation and cell subgroup identification simultaneously. A workflow of JRDNN-KM is presented in Figure 1. Significantly advancing from the published studies, by exploiting the neural network’s capacity for flexible function approximation, JRDNN-KM can capture various nonlinear dependencies and interactions among genes in different cell subgroups while concurrently handling zero-inflation. By integrating sparse and similarity regularization, and through the inclusion of both homogeneous and heterogeneous neurons in the hidden layers, we ensure that the inferred networks capture both shared and subgroup-specific structures. This enables a more comprehensive joint analysis across distinct cell subgroups. The inferred networks are further integrated into the Mahalanobis distance-based K-means clustering procedure to identify previously unknown cell subgroups. This approach offers the distinct advantage of simultaneously accounting for heterogeneity in both expression levels and network connectivity between genes. We demonstrate the performance of JRDNN-KM through extensive simulations and real single-cell transcriptomic datasets from humans and mice. JRDNN-KM detects biologically sensible cell subgroups, hub genes, and modules, significantly contributing to revealing the biological mechanisms driving cellular processes.
@@
Figure 1: Workflow of JRDNN-KM. (A) Input: normalized single-cell transcriptome data $\boldsymbol{x}_{i}$’s and initialized cell subgroup memberships $C_{i}$’s. (B) JRDNN-KM: iterations between JRDNN and Mahalanobis distance-based K-means clustering. (C) Output: $K$ networks for the $K$ subgroups, with both common and specific edges, and estimated cell subgroup memberships. (D) JRDNN architecture: The architecture comprises a selection layer and hidden layers with combined homogeneous and heterogeneous neurons, and is optimized by a zero-inflated loss function that includes sparse and similarity regularization terms.
2 Data Description
We analyze five technically diverse single-cell transcriptomic datasets from human and mouse systems, spanning a wide spectrum of tissue types, developmental stages, and physiological contexts. Specifically, these datasets include:
  • A human lung adenocarcinoma (LUAD) cell line (Tian et al., 2019), widely used in cancer research to investigate tumor heterogeneity;
  • Human peripheral blood mononuclear cells (PBMC) (Zheng et al., 2017), a well-characterized immune cell reference with direct relevance to immunology and biomarker discovery;
  • Mouse embryonic stem cells (mESCs) (Kolodziejczyk et al., 2015), frequently employed in studies of early development and pluripotency;
  • Mouse liver and uterus tissues from the Mouse Cell Atlas (MCA) (Han et al., 2018), two distinct datasets that capture tissue-specific complexity and organ-level functions.
These five datasets exhibit substantial variation in sample size, from the smaller mESC data to the large-scale PBMC data; data quality, ranging from profiles with low zero-inflation (mESCs) to those with high dropout rates exceeding 90% (PBMC); and subgroup distribution, including notably imbalanced cell type compositions as seen in the mouse uterus data. Each dataset is accompanied by carefully annotated cell identities based on established markers, serving as gold standards for evaluating cell subgroup detection performance. Additional details are provided in Supplementary Section S1.
Building on these rich and varied data resources, we aim to develop a novel statistical framework for heterogeneous network estimation that addresses the following two core scientific problems:
  1. (Q1)
    How can we reliably identify cell subgroups defined by distinct patterns of gene-gene relationships, while handling the high sparsity and technical noise (e.g., dropout events) inherent in single-cell transcriptomic data?
  2. (Q2)
    How can we model complex, nonlinear gene-gene relationships to reconstruct accurate and subgroup-specific gene networks and effectively explore the similarities and differences among these networks across different cell subgroups?
The simultaneous conduct of network analysis and cell subgroup identification addresses the fundamental need to uncover previously unrecognized heterogeneity and to reveal interrelationships among molecular features across different subgroups. This practical goal directly motivates our first question (Q1), which focuses on identifying cell subgroups characterized by distinct gene-gene relationships while effectively handling technical noise. Furthermore, as articulated in the second question (Q2), moving beyond linear assumptions to model complex, nonlinear gene-gene relationships enables a more accurate representation of molecular relationships and helps to overcome the limitations of conventional Gaussian graphical models. (Q2) is further motivated by the recognition that the analysis of cellular heterogeneity should be complemented by an examination of underlying homogeneity across subgroups, particularly when cells originate from the same tissue. Such integration is essential for identifying molecular programs shared across subgroups and network alterations unique to each subgroup, thereby providing a systemic view of gene relationship modulation in different biological contexts. Together, addressing these questions establishes a powerful framework for deciphering the molecular mechanisms that underlie key processes such as development and disease pathogenesis.
3 Proposed JRDNN-KM Method
@@
Suppose there are $n$ independent cells from $K$ subgroups. For the $i$th cell, denote $C_{i}\in\{1,\ldots,K\}$ as the subgroup assignment and $\boldsymbol{y}_{i}=\left(y_{i1},\ldots,y_{ip}\right)^{\top}$ as the $p$-dimensional vector of the observed count transcriptomic data. In practice, $\boldsymbol{y}_{i}$ has a high sparsity with zero-inflation due to the dropout event. Following the published studies (Booeshaghi and Pachter, 2021; Hafemeister and Satija, 2019), we conduct normalization using the R package Seurat and $\log(1+x)$ transformation on $\boldsymbol{y}_{i}$’s to accommodate the library size and count nature and denote the processed continuous vector as $\boldsymbol{x}_{i}=\left(x_{i1},\ldots,x_{ip}\right)^{\top}$.
@@ 3.1 Network estimation based on a joint regularized deep neural network
First, assume that the subgroup assignment is known, and for each subgroup, each gene is standardized to have a zero mean and unit variance. For notational simplicity, we preserve $\boldsymbol{x}_{i}$ to represent the standardized data. To tackle the challenges posed by the high sparsity of single-cell transcriptomic data (Q1) and the need to model nonlinear relationships in gene networks (Q2), when $C_{i}=k$, we propose a zero-inflated conditional Gaussian distribution as:
$x_{ij}|\boldsymbol{x}_{i,\backslash j};C_{i}=k\sim$ $\mathcal{N}\left(f_{kj}(\boldsymbol{w}_{k,j,\backslash j}*\boldsymbol{x}_{i,\backslash j}),\sigma_{kj}^{2}\right)\pi_{j}+\delta_{0}(x_{ij})(1-\pi_{j}),$ (1)
[…]
@@
[…]
@@
[…]
@@
with $l_{n}(\boldsymbol{x}_{i};\boldsymbol{U}_{kj},\boldsymbol{w}_{k,j,\backslash j},\pi_{j})=\delta_{ij}\left[\log\pi_{j}-\frac{1}{2}\left\{\log(2\pi)+\log\left(\sigma_{kj}^{2}\right)+\frac{\left(x_{ij}-g(\boldsymbol{x}_{i,\backslash j};\boldsymbol{U}_{kj},\boldsymbol{w}_{k,j,\backslash j})\right)^{2}}{\sigma_{kj}^{2}}\right\}\right]+(1-\delta_{ij})\log(1-\pi_{j}),$ where $\boldsymbol{X}$ is a $n\times p$ matrix consisting of $\boldsymbol{x}_{1},\cdots,\boldsymbol{x}_{n}$, $\Phi$ denotes all unknown parameters, $1(\cdot)$ is an indicator function, $\tau_{ik}=1(C_{i}=k)$, $\delta_{ij}=1(x_{ij}\neq 0)$, and $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ are three non-negative tuning parameters.
In (3.1), the first term is the negative log-likelihood function, which models the fit of our data. The second term, an $\ell_{1}$ penalty, encourages sparsity in the selection layer parameters $w_{k,jl}$’s to facilitate sparse network estimation. The third and fourth terms model the commonality among cell subgroups, thereby addressing the need to incorporate inter-subgroup network similarities and differences in (Q2). Specifically, for any pairs of $j$ and $l$, the third term promotes $1(w_{k,jl}\neq 0)=1(w_{k^{\prime},jl}\neq 0)$, resulting in that $w_{k,jl}$ and $w_{k^{\prime},jl}$ tend to be zero or nonzero simultaneously. Thus, the $K$ subgroups potentially share some common edges. We further introduce the fourth term to accommodate the similarity in connection strength of the edges. Specifically, the fourth term imposes the F-norm on the weights $\boldsymbol{\Delta}^{(1)}_{k}$ of heterogeneous neurons in the first layer, promoting the elements of $\boldsymbol{\Delta}^{(1)}_{k}$ shrunk towards zero simultaneously and thus pruning all heterogeneous neurons in the $M$ hidden layers. With (3.1), the constructed networks of different cell subgroups not only reflect heterogeneity but also exhibit common connections with similar strength.